OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 3x. What is the value of (f + g)(-1)?(i tried and got-x-8 but i think its wrong)

OpenStudy (anonymous):

(f+g)(-1)=f(-1)+g(-1)=10-3=7